### v1

# STEPC: A Pixel-wise Nonuniformity Correction Framework for Photon-Counting CT in Multi-material Imaging Scenarios

Abstract
Photon-counting computed tomography (PCCT) has demonstrated significant advancements in recent years; however, pixel-wise detector response nonuniformity remains a key challenge, frequently manifesting as ring artifacts in reconstructed images. Existing correction methods exhibit limited generalizability in complex multi-material scenarios, such as contrast-enhanced imaging. This study introduces a Signal-to-Uniformity Error Polynomial Calibration (STEPC) framework to address this issue. STEPC first fits multi-energy projections using a 2D polynomial surface to generate ideal references, then applies a nonlinear multi-energy polynomial model to predict and correct pixel-wise nonuniformity errors. The model is calibrated using homogeneous slab phantoms of different materials, including PMMA, aluminum, and iodinated contrast agents, enabling correction for both non-contrast and contrast-enhanced imaging. Experiments were performed on a custom Micro-PCCT system with phantoms and mouse. Correction performance of STEPC was evaluated using the mean local standard deviation (MLSD) in the projection domain and the ring artifact deviation (RAD) on the reconstructed images. STEPC consistently outperformed existing correction methods in both non-contrast and contrast-enhanced scenarios. It achieved the lowest MLSD and RAD for both phantoms and mouse scans. These results indicate that STEPC provides a robust and practical solution for correcting detector nonuniformity in multi-material PCCT imaging, witch position it as a promising general-purpose calibration framework for photon-counting CT systems.
{IEEEkeywords}
photon counting CT, detector nonuniformity correction, polynomial calibration, ring artifact suppression, contrast-enhanced imaging.
1 Introduction
[…]
Photon-counting computed tomography (PCCT) has gained considerable attention [1, 2] due to its advantages in ultra-high-resolution imaging [2], beam hardening reduction [3], material decomposition [4], K-edge imaging [5] and more. However, PCCT still faces several challenges, including energy threshold bias and nonuniform pixel responses arising from variations in detector crystal and post-processing electronics. These effects degrade energy resolution and cause pixel-level response nonuniformity, often resulting in ring artifacts in the reconstructed images [6, 7, 8, 9].
Various strategies have been proposed to suppress or correct pixel-wise nonuniformity in photon-counting CT, which can be broadly categorized into three main approaches: (1) Energy threshold calibration, which directly calibrates detector’s energy thresholds using monochromatic synchrotron sources, radioactive isotopes, X-ray fluorescence, or K-edge materials [10, 11]; (2) Ring artifact removal methods, which suppress artifacts in the projection or image domain through post-processing [12, 13, 14]; (3) Phantom-based measurement calibration, which uses calibration scans of known materials to learn a mapping from measured signals to ideal responses [15, 16, 17, 18]. Each approach has its own advantages and limitations. Energy threshold calibration can directly fix detector threshold bias but typically requires monochromatic X-ray sources and detector repositioning, witch is not readily available for most researchers. Ring artifact removal is simple to apply, but their correction capability is often limited, particularly in the presence of low-frequency concentric artifacts [7] or more severe nonlinear detector behaviors [19]. Phantom-based calibration directly models detector nonuniformity characteristics and is widely used as a key preprocessing step before reconstruction.
Among phantom-based measurement calibration methods, traditional single-energy correction techniques [20, 15, 21, 22] can improve uniformity in single-material scenarios but perform poorly when multiple materials (e.g., soft tissue, bone, or iodine contrast agents) are present. Material decomposition methods [23, 24, 17] could address multi-material calibration, but their practicality is restricted by current photon-counting detectors (PCDs) limited energy thresholds (typically $\leq 2$), allowing only dual-material decomposition. This renders them inadequate for complex tasks, such as contrast-enhanced imaging. Additionally, most methods demand precise phantom positioning and thickness measurements, increasing operational complexity.
To overcome these challenges, we propose Signal-to-Uniformity Error Polynomial Calibration (STEPC), a novel framework for predicting and correcting photon-counting detector uniformity error responses. The method consists of four main steps. First, ideal reference projections are generated by applying a second-order 2D polynomial fit to flat-field projections acquired from combinations of calibration materials, avoiding the need for precise phantom alignment and measurement of material thickness or density. Then, The nonuniformity error is calculated by comparing the ideal projections with the actual measured projections. Next, A multi-energy polynomial regression model is fitted to learn the nonlinear relationship between the measured multi-energy projections and the corresponding nonuniformity errors. Finally, during actual imaging, the trained model is used to predict and correct detector nonuniformity in the acquired projections.
Unlike material decomposition approaches, the proposed method is robust to multi-material objects even under limited energy threshold numbers. By combined multi-material calibration slab phantoms (e.g., PMMA, Al, and contrast agents), the framework adapts to various imaging scenarios, including both non-contrast and contrast-enhanced imaging. The key contributions of this work include:
(1) A novel calibration framework for correcting PCDs nonuniformity in multi-material imaging scenarios, which is easy to implement and remains effective even with only two energy thresholds.
(2) Design of multi-material calibration slab phantoms, including PMMA, AL, and iodixanol, enabling effective correction for both general and contrast-enhanced imaging.
(3) Experimental validation on both phantoms and mouse confirms the necessity of material-specific calibration and demonstrates the superior performance of the proposed method in multi-material imaging scenarios.
2 Related Works
Phantom-based measurement calibration methods typically involve three steps: (1) acquiring all possible incident spectra projections using various homogeneous slab phantoms, (2) calculating the ideal reference projections of the phantom, and (3) fitting a parametric model to map the measured signals to the reference ideal signals, which is then used for correcting actual object imaging. The main differences among these methods lie in the choice of incident spectra, and the form of the ideal signal and correction model.
Typically, different incident spectra are obtained using homogeneous phantoms with difference material (such as PMMA and AL) and thickness. The reference signal is commonly selected as either a single pixel value or the mean value [20, 16]. Alternative approaches include: (1) 1D polynomial fitting of projections to account for spatial variations in X-ray intensity or phantom thickness [25]. (2) directly using of measured thickness values as reference (common in material decomposition methods) [15, 18, 21, 22], and (3) simulation of ideal signals using known geometry, source spectra, and detector response functions [26, 27, 28].
Correction models are generally divided into traditional single-energy and multi-energy approaches. Single-energy methods perform energy-bin-wise corrections using linear [20], exponential [15], or empirical nonlinear functions [26]. While these methods work well for single-component objects, their correction accuracy degrades significantly for multi-material cases involving both soft tissue and bone. In contrast, multi-energy approaches leverage richer spectral information to enable more effective multi-material correction [16, 24, 17]. In the following, we introduce two representative single-energy methods: Flat-field Correction (FF) [29] and Signal-to-Equivalent Thickness Calibration (STC) [15], and two multi-energy methods: Affine Transformation Calibration (ATC) [16] and Polynomial Material Decomposition Calibration (PMDC) [24, 17], which will serve as comparative baselines in our study.
Figure 1: Distribution of spectral projection values for the non-contrast (a) and iodine-contrast (b) mouse scans.
(1) Flat-field Correction (FF)
Flat-field correction is the most common approach in energy-integrating CT systems. It includes dark-field (offset) and gain correction [29]. The corrected signal is computed as:
$I^{\mathrm{FF}}=\frac{N-B}{N_{\mathrm{air}}-B_{\mathrm{air}}}$ (1)
where $B$ and $B_{\mathrm{air}}$ are the dark-field signals for object and air scans, respectively, and $N$ and $N_{\mathrm{air}}$ are the raw counts of object and air scanning. In PCCT, the energy threshold can suppress electronic noise, $B\approx 0$ and $B_{\mathrm{air}}\approx 0$, yielding:
$I_{E}^{\mathrm{FF}}=\frac{N_{E}}{N_{E,\mathrm{air}}}$ (2)
where $E$ indexes the energy bin. The negative logarithmic transform is then applied for reconstruction:
$P_{E}^{\mathrm{FF}}=-\log(I_{E}^{\mathrm{FF}})$ (3)
(2) Signal-to-Equivalent Thickness Calibration (STC)
STC calibrates detector response using varying thicknesses of a single calibration material, typically PMMA for soft-tissue equivalence [15]. For each energy bin $E$, an exponential model is fit:
$N_{E}=C_{E}e^{A_{E}T_{E}}$ (4)
where $N_{E}$ is the detected photon count, $T_{E}$ is the material thickness, and $A_{E}$, $C_{E}$ are fitted parameters. The corrected equivalent thickness is calculated as:
$T_{E}^{\mathrm{STC}}=\frac{1}{A_{E}}\ln\left(\frac{N_{E}}{C_{E}}\right)$ (5)
This corrected thickness can be directly used for image reconstruction or converted into a virtual monoenergetic image for reconstruction.
(3) Affine Transformation Calibration (ATC)
Unlike FF and STC, which calibrate each energy bin data independently, ATC leverages all energy bin counts $N$ jointly to compensate for count loss caused by threshold variations [16]. The corrected counts are given by:
$\mathbf{N}^{\mathrm{ATC}}=\mathbf{A}\mathbf{N}+\mathbf{b}$ (6)
where $\mathbf{A}\in\mathbb{R}^{K\times K}$ and $\mathbf{b}\in\mathbb{R}^{K}$, with $K$ denoting the number of energy thresholds. The parameters $\mathbf{A}$ and $\mathbf{b}$ are learned by minimizing the mean squared error of the estimator for reference values $\lambda_{i}^{\mathrm{ref}}$:
$\mathrm{MSE}=\mathbb{E}\left[\left(\sum_{j=1}^{K}A_{ij}\widetilde{N}_{j}+b_{i}-\lambda_{i}^{\mathrm{ref}}\right)^{2}\right]$ (7)
where the expectation value is taken over all possible incident spectra and all possible noise realizations. The reference values $\lambda_{i}^{\mathrm{ref}}$ typically computed as the average value of projection.
(4) Polynomial Material Decomposition Calibration (PMDC)
Biological objects often consist of multiple materials, such as soft tissue and bone. The PMDC method corrects for multi-material effects by performing projection-domain material decomposition using polynomial fitting [24, 17]. Given flat-field corrected logarithmic projections $P_{E}=-\log(I_{E}^{\mathrm{FF}})$, the thicknesses of two basis materials are estimated as:
$T_{M_{1}}^{\mathrm{PMDC}}=\sum_{i+j\leq p}c_{ij}P_{E_{1}}^{i}P_{E_{2}}^{j}$ (8)
$T_{M_{2}}^{\mathrm{PMDC}}=\sum_{i+j\leq p}d_{ij}P_{E_{1}}^{i}P_{E_{2}}^{j}$ (9)
where $p$ is the polynomial order (typically 3), and the coefficients $c_{ij}$ and $d_{ij}$ are determined from calibration scans using difference thicknesses combinations of two basis materials.
Among the above methods, FF and STC calibrate each energy bin independently, ignoring the additional information between energy bins. Moreover, FF does not consider object-induced changes in spectra. STC accounts for spectral changes but assumes a single material, which limits accuracy for biological tissues composed of multiple components (e.g., soft tissue and bone). Some methods extend single-energy correction to multi-material scenarios, such as Feng et al.[7] estimated the thickness of one material, then segmented other components using soft thresholding, but requiring additional steps such as segmentation, forward, and backward projections. Other approaches, PETC[21] transforms aluminum-equivalent thickness into PMMA-equivalent thickness using bin-wise projection models, but requires iterative per-pixel Gauss-Newton optimization (up to 1000 iterations), resulting in high computational cost. ATC allows for varying incident spectra but assumes a linear detector response, which limits its ability to handle the nonlinear characteristics of practical systems. PMDC improves correction by incorporating multi-material spectral changes and multi-energy bin information, but remains limited by the number of energy thresholds. Most commercial PCCT systems offer only two thresholds, witch allows only dual-material decomposition and makes it unsuitable for contrast-enhanced imaging. As shown in Figure 1, the incident spectra of non-contrast mouse scans can be modeled using PMMA and aluminum, whereas iodine-enhanced cases require iodine as an additional calibration material. Moreover, PMDC relies on accurate phantom thickness measurements and involves a signal-to-thickness conversion, which introduces additional uncertainty and may amplify noise.
3 Methods
Figure 2: Workflow of the proposed STEPC method. Step 1: Estimation of ideal projections for calibration slab phantoms; Step 2: Calculation of pixel-wise nonuniformity errors; Step 3: Fitting of the empirical multi-energy polynomial calibration model; Step 4: Projection correction for various imaging scenarios (red arrows indicate noticeable pixel response nonuniformity before correction).
To help explain the principle of the proposed method, the physical origin of the non-uniform detector response in photon-counting detectors (PCDs) is first described in this section. For the $j$th pixel in energy bin $E_{k}$, the measured detector counts can be expressed as:
[…]
where $S(E)$ is the input X-ray spectrum, $E$ denotes the photon energy, $R_{j}(E^{\prime},E)$ represents the pixel-specific photon-counting detector response function, $E^{\prime}$ denotes the detected energy, $E_{\max}$ is the maximum photon energy determined by the tube voltage. Assuming the existence of an ideal or reference response function $R_{j}^{ideal}(E^{\prime},E)$ that ensures uniformity across pixels, the corresponding ideal detector counts is:
[…]
Thus, the nonuniformity response error is:
[…]
We propose a STEPC method that directly utilizes multi-energy projection information to predict and correct detector nonuniformity for multi-material imaging. The detailed workflow is illustrated in Figure 2 and consists of the following steps:
Step 1: Ideal projection estimation. STEPC first measure the transmission through different combinations of multiple material slab phantoms. The log-transformed flat-field corrected projection data are then computed as $P(E_{k})=-\log(I_{E}^{FF})$ to reduce the influence of X-ray source and detector instabilities. Next, a second-order polynomial fit is applied to the 2D projection image $(x,y)$ for each energy bin $E_{k}$ to obtain the ideal projection $P_{ideal}(x,y,E_{k})$:
$P_{ideal}(x,y,E_{k})=\sum_{i+j\leq 2}b_{ij}^{(k)}x^{i}y^{j}$ (13)
Here, $(x,y)$ are the pixel coordinates, and the coefficients $b_{ij}^{(k)}$ are obtained by minimizing the mean square error between $P_{E_{k}}$ and $P_{ideal}(E_{k})$. We use MATLAB’s regress function [30] to solve for these coefficients. As shown in Figure 2(Step1), this polynomial fitting compensates for spatial variations in X-ray intensity and angular-dependent transmission thickness in the slab phantom, without requiring precise geometric measurements, thereby simplifying the calibration process.
Step 2: Nonuniformity error calculation. For each pixel and energy channel, the residual between the measured projection and the ideal projection is computed to obtain the pixel-wise nonuniformity error:
$\Delta P_{ideal}(x,y,E_{k})=P(x,y,E_{k})-P_{ideal}(x,y,E_{k})$ (14)
As shown in Figure 2(Step2), the nonuniformity errors is dependent on the incident spectrum, which varies with material type and thickness. For single-material cases, single-energy projections can roughly estimate the error. However, for multi-materials, single-energy projections fail to capture the error accurately. In contrast, it illustrates that dual-energy information more effectively characterizes the error surface, allowing accurate modeling of errors caused by multi-materials.
Step 3: Empirical calibration model fitting. An empirical polynomial model is used to predict nonuniformity error. By incorporating all energy-channel projections as input, the model captures spectral variations associated with different materials, enabling more accurate and robust correction in multi-material imaging scenarios. The model is formulated as follows:
$\hat{\Delta P}(x,y,E_{k})=\sum_{|\boldsymbol{\alpha}|\leq p}c_{\boldsymbol{\alpha}}^{(k)}\prod_{n=1}^{N}P(x,y,E_{n})^{\alpha_{n}}$ (15)
where $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{N})$ is the multi-index vector, $|\boldsymbol{\alpha}|=\sum_{n=1}^{N}\alpha_{n}\leq p$ is the total polynomial degree, $N$ is the number of energy thresholds, and $c_{\boldsymbol{\alpha}}^{(k)}$ are the polynomial coefficients for energy channel $E_{k}$. These coefficients are determined by minimizing the mean square error of the estimator for ideal error $\Delta P_{ideal}$:
$MSE=\mathbb{E}\left[\left(\hat{\Delta P}(x,y,E_{k})-\Delta P_{ideal}(x,y,E_{k})\right)^{2}\right]$ (16)
The expectation is taken over calibration data collected from homogeneous phantoms composed of different materials (e.g., PMMA, Aluminum, Iodixanol) to cover a wide range of possible incident spectra. The coefficients $c_{\boldsymbol{\alpha}}^{(k)}$ can be solved using regression tools such as MATLAB’s regress function[30].
Step 4: Correction. Finally, real object projection data are corrected by directly inputting the raw multi-energy projections into the fitted polynomial model to estimate the nonuniformity error, which is then subtracted to yield the corrected projection:
$P_{corr}(x,y,E_{k})=P(x,y,E_{k})-\hat{\Delta P}(x,y,E_{k})$ (17)
Notably, STEPC avoids the need for accurate thickness or density measurements of calibration phantoms by using 2D polynomial fitting to generate ideal reference projections. In addition, despite having only two energy thresholds, STEPC can incorporate iodine-based calibration to capture spectral variations from contrast agents, soft tissue, and bone, enabling effective correction in complex imaging scenarios.
Figure 3: Micro-PCCT system and multi-material slab phantoms. (a) Micro-PCCT imaging system; (b) slab phantom placement setup; (c) three types of slab phantoms: PMMA, aluminum, and iodixanol; (d) schematic of iodixanol phantom container; (e) combinations of PMMA and aluminum slabs with varying thicknesses; (f) combinations of iodixanol solution slabs with PMMA and aluminum slabs.
4 Experiments
4.1 Micro Photon Counting CT and Calibration Materials
As shown in Figure 3(a), the Micro Photon Counting CT (Micro-PCCT) system used in our experiments was jointly developed by Hainan University and United Imaging Life Science Instrument (LSI, Wuhan, China). It adopts a translate-rotate architecture where the object remains stationary and the gantrol rotates, reducing motion artifacts during imaging. The X-ray source was operated at 80 kV and 200 µA, and a 0.5 mm aluminum filter was used to suppress low-energy photons. The photon counting detector has a resolution of $2063\times 505$ pixels (after cropping peripheral invalid pixels), each with a 100 $\mu$$\times$ 100 $\mu$m pixel size, and supports two independently adjustable energy thresholds. The detector thresholds were set to 15 keV and 30 keV. Each energy channel uses a 12-bit counter capable of recording up to 4096 photons per acquisition. Each acquisition thus provides three energy bins: Total (15–80 keV) and High (30–80 keV), with the Low (15–30 keV) bin derived by subtraction: $\mathrm{Low}=\mathrm{Total}-\mathrm{High}$.
Figure 4: multi-material cylindrical phantoms. (a) Actual pictures of cylindrical phantoms; (b) CaCl2 inserts with concentrations of 100, 200, 400, and 600 mg/mL insert in a PMMA holder; (c) Iodixanol and CaCl2 inserts (iodixanol: 20, 50, 100 mg/mL; CaCl2: 200, 400 mg/mL); (d) PMMA only; (e) 200 mg/mL CaCl2 (2mm thickness PMMA cylindrical holder); (f) 50 mg/mL iodixanol.
Three types of slab phantoms were designed, as shown in Figure 3(c), including PMMA slabs with thicknesses of 0, 5, 10, 15, 20, 30, and 40 mm; aluminum slabs with thicknesses of 0, 0.5, 1, 1.5, 2, 3, 4, 5, 6, and 8 mm; and iodixanol solution slabs with concentrations of 0, 5, 10, 15, 20, 30, 50, 70, 100, 150, and 250 mg/cm3, sealed in 2 mm thick PMMA containers with a 6 mm solution core (Figure 3(d)). Two imaging scenarios were considered. For the non-contrast scenario, calibration was performed using combinations of PMMA and aluminum slabs, as shown in Figure 3(e). For the iodine-enhanced scenario, the same PMMA and aluminum slab combinations were used, and additional calibration was performed by incorporating iodixanol solutions combined with PMMA and aluminum slabs (0,0, 0,1, 0,3, 10,0, 10,1, and 20,0), as illustrated in Figure 3(f). For each slab combination, 600 projection frames were acquired and averaged to reduce noise.
4.2 Phantom and Mouse Imaging
Cylindrical Phantom Scanning
To evaluate correction performance under different conditions, five cylindrical phantoms were scanned (Figure 4): PMMA-only cylinders, 200 mg/mL CaCl2 cylinders (PMMA container with 2 mm thickness), 50 mg/mL Iodixanol cylinder (PMMA container with 2 mm thickness), cylinders with inserts of CaCl2 at $\{100,200,400,600\}$ mg/mL, cylinders with inserts of CaCl2 $\{200,400\}$ mg/mL and iodixanol $\{20,50,100\}$ mg/mL. All phantoms had an outer diameter of 30 mm, and the inserts diameter is 8 mm. Projections were acquired in “Continuous” mode with 1440 views per rotation (0.25° per view), a field of view (FOV) of 50 mm, a source-to-isocenter distance (SID) of 90 mm, and a source-to-detector distance (SDD) of 325 mm. All other scan settings were the same as in the calibration step.
Mouse Scanning
To validate correction performance in vivo, two C57BL mice (10 weeks old, $\sim$23 g) were scanned. One mouse underwent a non-contrast head scan, while the other received an intravenous injection of 0.3 mL iodixanol (300 mg/ml) via the tail vein and was scanned in the abdominal kidney region. Both mice were euthanized through intraperitoneal anesthesia prior to scanning to eliminate motion artifacts. Imaging parameters were the same as those used for the cylindrical phantom, except for FOV = 35 mm, SID = 74 mm, SDD = 325 mm.
All animal experiments were conducted using mice maintained in SPF animal facilities. Animal care and experimental protocols were approved by the Institutional Animal Care and Use Committee (IACUC) and the Ethical Committee of Animal Experiments of the School of Biomedical Engineering at Hainan University (Approval number: HNUAUCC-2022-00091).
4.3 Data Processing and Reconstruction
Detector nonuniformity correction was performed in the projection domain. The detector comprises $16\times 2$ tiles, separated by gap pixels and affected by defective pixels. These bad pixels were pre-identified and corrected via linear interpolation. Image reconstruction was carried out using the open-source TIGRE toolbox [31], employing the FDK algorithm with Hann windowing, resulting in volumes of size $1529\times 1529\times 400$. Images were converted to Hounsfield Units (HU) using a water phantom for calibration. All processing was performed in MATLAB R2024a [32].
Figure 5: Illustration of the Ring Artifact Deviation (RAD) calculation. Left: Reconstructed image of a PMMA phantom shown in polar coordinates $(r,\theta)$. Right: Angularly averaged intensity profile. A second-order polynomial fit is used to generate the ideal baseline, which is subtracted from the real averaged profile to remove cupping artifacts.
4.4 Baselines
We compare our STEPC method with four baselines: FF, STC, ATC, and PMDC. Since these methods work in different domains, appropriate post-processing was applied to ensure a fair comparison:
  • FF: Logarithmic projections were directly computed as $P_{E}^{FF}=-\log(I_{E}^{FF})$.
  • STC: Projections were computed using the calibrated thickness model: $P_{E}^{STC}=-\log\left(\frac{N_{E}^{STC}}{\bar{N}_{E,\mathrm{air}}}\right)$, where $N_{E}^{STC}=\bar{C}_{E}e^{\bar{A}_{E}T_{E}}$ and $\bar{C}_{E}$, $\bar{A}_{E}$ are average calibration coefficients obtained across all pixels.
  • ATC: Instead of using mean counts as described in Ref. [16], the same second-order 2D polynomial fit was applied to generate reference photon counts. After ATC correction in the counting domain, logarithmic air normalization was performed: $P_{E}^{ATC}=-\log\left(\frac{N_{E}^{ATC}}{\acute{N}_{E,\mathrm{air}}}\right)$. $\acute{N}_{E,\mathrm{air}}$ was also obtained by applying a second-order 2D polynomial fit to the air scan projection.
  • PMDC: For each energy bin $E(k)$, the following polynomial forward model was fitted:
    $P_{E(k)}=\sum_{i+j\leq 3}c_{ij}^{(k)}(T_{\mathrm{PMMA}})^{i}(T_{\mathrm{Al}})^{j},$ (18)
    After the PMDC correction obtaining $T_{\mathrm{PMMA}}^{PMDC}$ and $T_{\mathrm{Al}}^{PMDC}$, averaged coefficients $\bar{c}_{ij}^{(k)}$ across all pixels were used to compute corrected projections:
    $P_{E(k)}^{PMDC}=\sum_{i+j\leq 3}\bar{c}_{ij}^{(k)}(T_{\mathrm{PMMA}}^{PMDC})^{i}(T_{\mathrm{Al}}^{PMDC})^{j}.$ (19)
Note: FF requires only air scans for calibration; STC used only PMMA slabs for calibration; ATC and STEPC used PMMA and aluminum slabs for the non-contrast scenario and additional iodixanol slabs for the contrast scenario; PMDC used only PMMA and aluminum for both scenarios due to its two-threshold limitation. The polynomial order for both PMDC and STEPC was set to three.
4.5 Metrics
To assess the proposed correction method, we calculated the Mean Local Standard Deviation (MLSD) using a $20\times 20$ pixels window across each corrected projections. Lower values indicate better projection uniformity after correction.
To assess ring artifact suppression in reconstructed images, we used the Ring Artifact Deviation (RAD) metric described by Rodesch et al. [11]. As shown in Figure 5, images were first converted to polar coordinates centered on the rotation axis. Angular averaging was then applied, followed by a second-order polynomial fit along the radial direction to remove cupping artifacts. Finally, the standard deviation of the residuals across valid radial ranges and all slices was calculated. Since the single material phantom and multi-material insert phantoms have different internal structures, we selected different ROI sizes when computing the RAD metric to ensure the regions uniformly, so that the measurement reflected only ring artifacts intensity. For the single-material cylindrical phantoms, a 20 mm diameter ROI was selected, while for the insert phantoms, an 8 mm ROI was used. For the mouse scans, due to the complexity of anatomical structures, smaller ROIs were chosen to ensure regional uniformity. Specifically, a 3.5 mm diameter ROI was used for the head scan, and a 3.7 mm diameter ROI was used for the kidney scan, as indicated by the red circles in Figure 11. However, the same ROI was applied across all methods for each dataset, ensuring fair and consistent comparisons.
5 Results
We evaluated the corrected projection uniformity for different material slabs, and ring artifact suppression in cylindrical phantoms and mouse images. First, we quantitatively assessed projection uniformity for all PMMA+Al slab combinations (Figures 7) and PMMA+Al+iodixanol phantom slabs (Figures 8), with averaged MLSD results shown in Table 1. Second, we compared ring artifact suppression qualitatively (Figures 91011) and quantitatively (Tables 2) for phantoms and mouse imaging. Finally, we performed a sensitivity analysis (Table 3, Figure 12) to evaluate the impact of polynomial degree and calibration material choice on difference material objects imaging.
5.1 Projection Uniformity for Calibration Slabs
We evaluated projection uniformity after applying different correction methods across various combinations of materials and thicknesses. Figure 6 presents the results for one representative case (10 mm PMMA + 1.5 mm Al). Among all methods, FF correction resulted in the poorest uniformity, while the proposed STEPC method achieved the most consistent and uniform projections.
Figure 6: Corrected projections for the 10 mm PMMA + 1.5 mm Al case using different methods: (a) FF, (b) STC, (c) ATC, (d) PMDC, (e) STEPC, and (f) the corresponding projection profiles.
Figure 7: Non-contrast scenario calibration: mean local standard deviation maps of corrected projections for PMMA and aluminum slab combinations using different correction methods: (a) Low, (b) High, (c) Total.
Figure 8: Iodine-enhanced scenario calibration: mean local standard deviation maps of corrected projections for PMMA, aluminum and iodixanol slab combinations using different correction methods: (a) Low of PMMA+AL, (b) High of PMMA+AL, (c) Total of PMMA+AL, (d) Low of PMMA+AL+Iodixanol, (e) High of PMMA+AL+Iodixanol, (f) Total of PMMA+AL+Iodixanol.
For the Non-contrast scenario calibration, the projection MLSD maps for all PMMA and Al combinations are shown in Figures 7. FF consistently performed the worst across all cases. STC performed well with PMMA alone but quickly deteriorated when aluminum was added due to the lack of aluminum calibration. ATC performed well for thin materials but degraded with increasing thickness, especially in the Low and Total energy bins, due to stronger nonlinear effects at lower photon energies. PMDC performed poorly with PMMA-only phantoms, likely due to instability in the material decomposition process. In contrast, STEPC consistently achieved near-best uniformity under all conditions.
For the iodine-enhanced scenario calibration, projection MLSD maps for all PMMA, aluminum, and iodixanol combinations are shown in Figures 8. For the only PMMA and Al combinations as showed in Figures 8(a-c), the results of FF, STC, and PMDC remain unchanged, as they did not incorporate iodixanol in calibration. Only ATC and STEPC added iodixanol in calibration, resulting in updated results. ATC showed increased MLSD for PMMA+Al projections due to the added material, while STEPC maintained stable performance, with only a slight increase in the PMMA-only projections. For combinations including iodixanol as showed in Figures 8(d-f), FF still showed the poorest performance. STC also failed to correct projections containing both aluminum and iodixanol. PMDC’s performance degraded sharply with higher iodixanol concentrations. In contrast, ATC performed better overall with the inclusion of all three materials. STEPC again achieved the most uniform projections across all material combinations. Table 1 summarizes the average MLSD values, showing that STEPC consistently yielded the lowest MLSD across both iodine-free and iodine-enhanced scenarios in all energy bins.
Table 1: Mean local standard deviation (MLSD) of different methods
Calibration Materials Energy MLSD(×$\times 10^{-2}$$\downarrow$
FF STC ATC PMDC STEPC
PMMA + AL Low 5.09 1.58 1.31 0.50 0.38
High 2.28 1.10 0.31 0.19 0.18
Total 1.40 0.41 0.26 0.24 0.19
PMMA + AL + Iodixanol Low 5.31 1.09 0.90 1.81 0.56
High 1.13 0.76 0.35 0.75 0.25
Total 1.34 0.29 0.24 0.82 0.21
5.2 Cylindrical Phantom Imageing
Figure 9 shows reconstructed images of cylindrical phantoms without iodine contrast. For the PMMA cylinder, FF exhibited severe rings, while ATC and PMDC showed mild artifacts, and STC and STEPC produced the least ring artifacts. For the CaCl2 and CaCl2 insert phantom, STEPC and PMDC delivered the best performance, whereas STC and ATC still showed visible artifacts, especially in low-energy images. These observations align with the earlier slab calibration results: STC was calibrated using only PMMA, so it performed well for PMMA but poorly for other materials. PMDC, on the other hand, showed residual artifacts in the PMMA phantom due to decomposition instability when applied to single-material objects. In contrast, STEPC maintained robust performance across both single- and multi-material cases.
Figure 9: Reconstructed images of the phantom without contrast using different correction methods. Display window: [-1000, 3000] HU.
Figure 10 shows reconstructed images of cylindrical phantoms containing iodine contrast. For the iodixanol-only phantom, FF exhibits severe ring artifacts, particularly in the low-energy bin. STC still shows noticeable artifacts, while ATC leaves only mild residuals. PMDC, however, suffers from prominent ring artifacts. In contrast, STEPC effectively suppresses ring artifacts. For the multi-material inserts (CaCl2 + iodixanol), similar trends are observed. PMDC continues to produce strong artifacts due to the influence of iodine inserts, highlighting its limitation in contrast-enhanced scenarios. Although ATC performs better at the center, visible artifacts remain between the inserts, as indicated by the red arrows. STEPC consistently produces artifact-free images across the entire field, demonstrating its robustness for multi-material correction.
Figure 10: Reconstructed images of the phantom with iodine contrast using different correction methods. Display window: [-1000, 3000] HU.
Table 2 quantifies the RAD for all phantoms, consistent with the previous qualitative observations. STEPC achieves the lowest RAD in nearly all cases, except for the high-energy PMMA images, where STC performs best. This is likely because STC was calibrated using only PMMA, and PMMA exhibits better linear attenuation at high energie bin, which better matches the linear attenuation assumptions inherent in the STC model.
Table 2: Ring artifact deviation (RAD) of different methods on phantom and mouse images
Objects Energy RAD (HU) ↓
FF STC ATC PMDC STEPC
No Contrast Phantoms PMMA Low 279.19 27.07 87.87 47.09 25.99
High 135.20 35.48 52.13 45.29 35.75
Total 122.26 17.12 29.25 39.17 16.93
200mg/ml CaCI2 Low 389.03 196.94 168.67 50.43 39.39
High 419.77 212.37 50.07 34.75 31.86
Total 170.27 46.70 30.50 37.16 21.75
CaCI2 inserts Low 478.42 61.65 80.27 47.46 47.08
High 285.94 106.85 55.82 46.58 44.49
Total 196.21 36.02 34.74 33.08 31.12
Contrast Phantoms 50mg/ml Iodixanol Low 500.56 69.76 66.96 375.98 34.85
High 116.66 174.31 54.08 272.72 39.80
Total 167.53 56.71 23.78 218.23 21.99
CaCI2 + Iodixanol inserts Low 474.48 51.96 39.72 87.50 34.45
High 221.68 55.61 43.07 65.00 35.66
Total 193.17 30.19 27.03 56.46 24.49
Mouse Head (without contrast) Low 775.55 194.26 164.72 76.67 62.94
High 373.29 167.71 103.98 89.36 83.49
Total 339.10 94.25 56.88 56.45 40.98
Kidney (iodixanol contrast) Low 1116.58 140.50 131.03 283.79 86.19
High 425.00 177.35 115.76 234.82 102.59
Total 429.04 83.26 56.13 162.80 54.91
5.3 Mouse Imaging
Figure 11 shows reconstructed images of mouse scans, and Table 2 includes the corresponding quantitative RAD results. In the non-contrast mouse head, FF produced severe ring artifacts that obscured anatomical details. STC also exhibited visible artifacts, while ATC and PMDC showed relatively milder ring artifacts. However, ATC displayed more pronounced artifacts in the low-energy bin, likely due to its limited ability to correct for the stronger nonlinear effects at lower energies. In contrast, STEPC achieved nearly complete artifact suppression and yielded the lowest RAD values across all energy bins. In the contrast-enhanced kidney images, FF and STC continued to show noticeable artifacts. Both STEPC and ATC performed well due to iodixanol calibration, but ATC showed slight artifacts in the low-energy bin, again likely due to nonlinear spectral effects. PMDC, however, displayed pronounced ring artifacts at the center and severe streak artifacts at the kidney edges, especially in low and total energy images (indicated by red arrows), highlighting its limitation in contrast-enhanced scenarios. STEPC consistently achieved the lowest RAD scores across all energy bins. Overall, STEPC consistently delivered the best performance in both non-contrast and contrast-enhanced mouse imaging.
Figure 11: Reconstructed mouse images of the non-contrast head and contrast-enhanced kidney. Display window: [-1000, 3000] HU.
5.4 Sensitivity Analysis
We evaluated the sensitivity of STEPC to polynomial order and calibration material selection, as shown in Figure 12 and Table 3. When using only PMMA and aluminum for calibration (labeled “w/o I”), RAD for PMMA and CaCl2 phantoms dropped quickly at orders below 2 but showed little improvement or even slight degradation at order 3, such as for PMMA at high energy. For the iodixanol phantom, RAD decreased up to second order but increased at third order, likely due to overfitting to PMMA and aluminum, reducing generalization to new materials. Therefore, a second-order polynomial is typically sufficient.
Table 3: Quantitative evaluation of STEPC with respect to polynomial order and calibration materials.
Setting Ring Artifact Deviation (HU) ↓
PMMA 200mg/ml CaCI2 50mg/ml Iodixanol
Low High Total Low High Total Low High Total
FF 279.19 135.20 122.26 389.03 419.77 170.27 500.56 116.66 167.53
1-order w/o I 119.40 40.55 52.88 114.80 37.32 41.47 304.87 123.63 74.50
2-order w/o I 28.27 32.81 17.86 39.19 31.16 22.12 176.35 117.92 46.91
3-order w/o I 25.99 35.75 16.93 39.39 31.86 21.75 263.50 202.38 103.85
3-order with I 59.98 45.56 20.31 38.87 31.63 21.63 34.85 39.80 21.99
When iodixanol slabs were included in calibration (labeled “with I”), RAD for iodixanol phantoms decreased significantly, indicating improved correction. However, RAD for PMMA slightly increased, and CaCl2 remained stable or improved slightly at low and total energies. This is because adding iodixanol forces the model to learn a more balanced representation across all three materials. Since iodixanol and PMMA have similar attenuation distribution and differ more from aluminum (see Figure 1), iodixanol reduces the model’s specificity to PMMA. However, this impact is limited in practice. Firstly, real biological imaging typically involves multiple materials (e.g., soft tissue, bone, and contrast agents), where the correction remains stable and effective for multi-component objects. More importantly, we typically know whether the contrast agent is present. For non-contrast cases, a model calibrated only with PMMA and aluminum can be used; for contrast-enhanced scans, adding contrast agents into the calibration process can significantly improve correction performance. In summary, a second-order polynomial is sufficient for STEPC, and including contrast materials in calibration is beneficial for contrast-enhanced imaging.
Figure 12: Sensitivity analysis of STEPC with respect to polynomial order and calibration materials.
6 DISCUSSION
In this study, we proposed the Signal-to-Uniformity Error Polynomial Calibration (STEPC) method to address detector response nonuniformity in photon-counting CT systems. It begins by generating ideal flat-field projections using a 2D second-order polynomial fit, eliminating the need for precise physical thickness measurements of calibration phantoms. An empirical polynomial model is then built on the residuals across all energy thresholds. This approach overcomes the limitation of requiring more energy thresholds for multi-material calibration, enabling accurate prediction and correction of nonuniformity errors under complex incident spectra.
Detector nonuniformity is dependent on the incident X-ray spectra as indicated in equation 12. Previous studies have used slab phantoms with varying thicknesses of PMMA (to simulate soft tissue) and aluminum (to simulate bone) to perform single or dual-material calibration. However, no calibration strategies have yet been developed specifically for contrast agents. We found that the choice of calibration materials is critical. As shown in Figure 9, STC calibrated only with PMMA performs well for PMMA but poorly for CaCl2 phantom. Similarly, PMDC, calibrated with both PMMA and aluminum, achieves better correction for both PMMA, CaCl2 phantoms and mouse head, but its performance degrades in the presence of contrast agents, as seen in Figure 10 and Figure 11. Notably, nonlinear spectral effects are more prominent at lower energy bins due to increased beam hardening and material-specific absorption variations. This significantly challenges calibration models that assume linearity, such as ATC, which exhibits more severe ring artifacts in low-energy images, as shown in Figure 9-11. To address these challenges, we designed a dedicated slab phantom containing an iodinated contrast solution to calibrate contrast-enhanced objects. However, most photon-counting detectors are currently limited to two energy thresholds, making material decomposition methods like PMDC unsuitable for three-material decomposition. The proposed STEPC method overcomes this limitation by directly modeling and predicting nonuniformity errors without relying on material decomposition and additional energy thresholds. By employing a nonlinear multi-energy polynomial model, STEPC effectively captures spectral nonlinearity and enables more accurate and robust correction in complex multi-material imaging scenarios.
On the other hand, the proposed STEPC method can be flexibly adapted to different imaging scenarios. For non-contrast-enhanced mouse scans, using only PMMA and aluminum in calibration is sufficient to approximate the possible incident spectra, as shown in Figure 1. For contrast-enhanced scans, only an additional slab phantom containing the contrast agent needs to be included, no changes to the model itself are required. STEPC also demonstrates strong robustness, successfully suppressing ring artifacts across single-material, dual-material, and even triple-material phantoms with iodinated contrast, as shown in Figure 9-10. In vivo results also confirm its superior performance in both non-contrast mouse head and contrast-enhanced kidney imaging (Figure 11). Moreover, STEPC offers greater implementation flexibility and low computational cost. Unlike thickness-based correction methods such as STC and PMDC, which require precise phantom positioning and accurate thickness measurements, STEPC uses a 2D second-order polynomial to estimate the ideal projection surface, simplifying the calibration process. Meanwhile, As demonstrated in the Figure 12 and Table 3, a second-order polynomial is sufficient to achieve high correction accuracy, significantly reducing computational demands.
Despite its advantages, this work has several limitations and directions for future research:
  1. 1.
    Optimization of Calibration Materials: The current use of various thickness or density combinations increases calibration complexity. Future work could optimize and reduce the number of combinations to simplify the calibration process.
  2. 2.
    Extension to Complex Scenarios: This study focuses on dual-threshold detectors. As future photon-counting CT systems support more energy thresholds, STEPC could be extended to higher-dimensional spectral spaces to better handle multi-contrast-agent cases. It is also necessary to assess the generalizability of iodine-based calibration to other contrast agents (e.g., gadolinium or barium).
  3. 3.
    Neural Network Models: STEPC uses empirical polynomial model for errors prediction. While efficient and interpretable, it may underperform in more complex scenarios involving multiple contrast agents or high noise levels. Future work could incorporate neural networks to improve correction accuracy and generalizability.
7 Conclusion
In this study, we proposed the STEPC method to address detector nonuniformity in photon-counting CT. STEPC enables effective correction for multi-material objects even with only two energy thresholds. Experimental results demonstrate that, compared to existing single- and multi-energy correction methods, STEPC more effectively leverages spectral information and shows superior robustness and generalizability, particularly in contrast-enhanced imaging, where it achieves better artifact suppression. Overall, STEPC offers greater operational flexibility and adaptability for complex multi-material scenarios. Therefore, these qualities position it as a promising general-purpose calibration framework for photon-counting CT systems.
[…]